A broken leg with a mass of 5kg receives an equivalent dose of 0.5mSv. If the X-ray energy is 50keV,how many x-ray photons were absorbed?


1
Expert's answer
2021-03-01T12:38:07-0500

DT.R = "\frac{H_T}{w_R}=\frac{0.5\times10^{-3}}{1} = 0.5\times10^{-3} Gy"

"\\epsilon_T = D_Tm_T = (0.5\\times10^{-3}Gy)(5\u00a0kg) = 2.5\\times10^{-3} J"

N = "\frac{\epsilon_T}{E_{photon}} = \frac{2.5\times10^{-3}J}{(50\times10^{-3}MeV)(1.602\times10^{-13}\frac{J}{MeV})}= 3.12\times 10^{11}"

Answer: N = 3.12"\times"1011.